Two times the least of three consecutive odd integers exceeds three times the greatest by 15. What are the integers?
Answer by CubeyThePenguin(3113)

consecutive odd integers: (x-2), x, (x+2)

2(x-2) = 3(x+2) + 15
2x-4 = 3x + 21
-25 = x

The integers are -27, -25, and -23.